OpenStudy (anonymous):

identify the slope and y-intercept of the line with the given equation. y=-7+5x

OpenStudy (anonymous):

5 is the slope y-intercept is -7

OpenStudy (anonymous):

equation of linear function is y = mx + b where m is the slope and b is the y-intercept

OpenStudy (anonymous):

y= mx+b slope= 5 y-intercept=-7

jimthompson5910 (jim_thompson5910):

y=-7+5x y=5x-7 The equation is now in slope intercept form y = mx+b where m is the slope and (0, b) is the y-intercept Since m = 5 and b = -7, the slope is 5 and the y-intercept is (0, -7)